OpenStudy (anonymous):

Scott earns $5 for each lawn he mows. On Saturday he mowed 6 lawns, and on Sunday he mowed 8 lawns. Which of the following expressions represents how much Scott earned for the weekend (in dollars)?

OpenStudy (uri):

total laws mowed =8+6 =14 Earning on each lawn mowed =5 Earned for the weekend=5*14 =70. @xmeowy